# Instability of a fluctuating biomimetic membrane driven by an applied uniform DC electric field

### v1

@@ III Linear stability analysis
The membrane fluctuations impose small perturbation $h(x,t)=h_{k}e^{st+ikx}$ around a flat reference state. Any variable $g$ is expanded in a series of the form $g=g^{(0)}(z,t)+g^{(1)}(z,x,t)$. The superscript $(0)$ corresponds to the base state, and $g^{(1)}=g_{k}^{(1)}e^{st+ikx}$ is the first order solution due to the effect of the small undulations. For the convenience, we will drop the subscript $k$ for the first order perturbation.
@@ III.1 Base state
We consider the flat membrane as the base state, $h(x)=0$, and the electric field is aligned in $z$ direction, perpendicular to the membrane. For electrostatic fields, we solve Eq. (13a) and Eq. (13b), with the Robin BC (15) and non-flux condition (7) at the membrane, and $\rho_{i}(z\rightarrow\pm\infty)=0$, $\phi_{i}(z\rightarrow\pm\infty)=\pm V$ at the far field, to obtain the charge distribution
$\rho_{1}^{(0)}=\rho_{m}e^{-\frac{z}{\gamma}},\qquad\rho_{2}^{(0)}=-\gamma^{2}\rho_{m}e^{\gamma z}$ (21)
and the base state potential
@@
is related to the jump in the charge density across the membrane, with ${[\![}\rho^{(0)}{]\!]}_{z=0}=(1+\gamma^{2})\rho_{m}$. Note that the base state current $\bm{j}_{1}^{(0)}=\bm{j}_{2}^{(0)}=\bm{0}$ for all $z$, not just at the membrane [24].
The base state field can be calculated as $\bm{E}_{i}^{(0)}=E_{z,i}^{(0)}\hat{e}_{z}=-\partial_{z}\phi_{i}^{(0)}\hat{e}_{z}$, which is non zero in the Debye layer. This represents a key difference from the base state field in LDM. In LDM, under the same physical setup, the base state field is zero unless there is ion current across the membrane [10, 7] or membrane capacitor is charging [12]. In our subsequent discussion, we will demonstrate the critical role played by the non-zero base state electric field $\bm{E}_{i}^{(0)}$ in the membrane instability. The transmembrane potential at base state is given by
@@
where the constant internal field has been used due to the continuity of the potential at the membrane boundaries, which constitutes with $\varepsilon_{m}E_{m}^{(0)}=\varepsilon_{f}E_{z,i}^{(0)}|_{z=0}$ by substituting $\beta$ in Eq. (16). The internal field $E_{m}^{(0)}$ is obtained by keeping a finite thickness $d$, and in Sec. III.3 we will show this field introduces negative tension and results into destabilization effect.
@@
by assuming $p_{i}^{(0)}(z\rightarrow\pm\infty)=0$. This indicates that the pressure is continuous ${[\![}p^{(0)}{]\!]}_{z=0}=0$ across the flat membrane.
III.2 Leading order equations
@@
which can be solved by the first order potential
@@
$a_{1}=\left[\frac{(\gamma^{2}+1)\beta+l_{2}}{l_{1}l_{2}+\beta(l_{1}+l_{2})}\right],\quad a_{2}=\left[\frac{(\gamma^{2}+1)\beta+\gamma^{2}l_{1}}{l_{1}l_{2}+\beta(l_{1}+l_{2})}\right].$ (31)
can be obtained via imposing the first order of Robin BC at the membrane
$\left(h\partial_{zz}\phi_{i}^{(0)}+\partial_{z}\phi_{i}^{(1)}\right)_{z=0}=\beta{[\![}\phi^{(1)}+h\partial_{z}\phi^{(0)}{]\!]}_{z=0},$ (32)
Note that when subjected to perturbation, the first order current remains zero in both electrolytes $\bm{j}_{1}^{(1)}=\bm{j}_{2}^{(1)}=\bm{0}$.
@@
Rearranging Eqs. (33) gives rise to a fourth-order differential equation in $v_{i}$ which is given by
@@
By imposing $v_{i}(z\rightarrow\pm\infty)=0$, we obtain the solution of the form $v_{i}^{(1)}=(B_{i}+C_{i}z)e^{\mp kz}$, where $B_{i}$ and $C_{i}$ are to be determined by appropriate boundary conditions in terms of perturbation variables. The continuity of the tangential velocity, together with the incompressibility of the interface at the first order impose $u_{1}^{(1)}|_{z=0}=u_{2}^{(1)}|_{z=0}=0$. This reduces the kinetic equation to be simplify $v^{(1)}=\partial_{t}h=sh$, together with the continuity of the normal velocity we will have $v_{1}^{(1)}|_{z=0}=v_{2}^{(1)}|_{z=0}=sh$. These BCs determine the solutions as function of $s,h$. The pressure $p^{(1)}_{i}=(\partial_{z}^{3}/{k}^{2}-\partial_{z})v_{i}^{(1)}-(\rho^{(0)}\phi^{(1)})$ can be obtained as
$p_{1}^{(1)}=2kshe^{-kz}-\rho_{1}^{(0)}\phi_{1}^{(1)},\qquad p_{2}^{(1)}=-2kshe^{kz}-\rho_{2}^{(0)}\phi_{2}^{(1)}.$ (35)
To determine the growth rate $s$, we need the normal traction balance at the interface
@@
The term $h\partial_{z}\left[-p^{(0)}+\frac{1}{2}(\partial_{z}\phi^{(0)})^{2}\right]$ vanishes due to the stress balance at the base state, and $\partial_{z}v^{(1)}=0$ due to the incompressibility of the membrane and the bulk fluids. From Eq. (36) we see three sources of electrostatic contribution to the membrane instability, i.e., one from the pressure difference ${[\![}-p{]\!]}$ due to the body force, one from the external normal traction ${[\![}\partial_{z}\phi^{(0)}\partial_{z}\phi^{(1)}{]\!]}$, and one from the internal normal traction ${[\![}-\partial_{z}\phi_{m}^{(0)}\partial_{z}\phi_{m}^{(1)}{]\!]}$ of the membrane. Note that while the base state field $-\partial_{z}\phi_{m}^{(0)}$ is constant, the first order field $-\partial_{z}\phi_{m}^{(1)}$ varies across the membrane (see Appendix A and [20] for detailed calculation). For a zero-thickness model, this internal traction is needed to correct the electrohydrodynamic effect as seen in [20] and [18].
@@ III.3 Linear growth rate
The growth rate can be determined from Eq. (36) as
@@
where $s_{\text{ex}}$ represents the external contribution to the growth rate, including both body forces and normal traction, whereas the internal contribution $s_{\text{in}}$ arises solely from normal traction:
$s_{\text{ex}}$ $=\frac{\rho_{m}^{2}}{k}[\underbrace{-(\gamma^{2}a_{2}+a_{1})}_{\text{external body force}}+\underbrace{\gamma(\gamma^{2}+1)}_{\text{external normal traction}}],$ (38a)
@@
The three components, along with the full growth rate in Eq. (37) is shown in Fig. 2. Note that in the zero thickness limit $d\rightarrow 0$, the internal normal traction term in Eq. (38b) reduces to
@@
contributing linearly to the growth rate. This term is in agreement with  [7] in the LDM framework.
Figure 2: The components of growth rate in Eq. (37) varing as a function of perturbation wavenumber $k$, with the “total” representing the full result as the sum of the external, internal, and elastic contributions. We consider symmetric conductivity with $\gamma=1$, $V=2$, $\beta=10$, $Ca=2$, $\bar{\Gamma}=0$ and $d=0$.
Next, we examine the effect of each terms in the long-wave limit $k\rightarrow 0$. With a finite thickness $d$, the internal normal traction can be expanded as
@@
The body force in Eq. (38a) can be expand as
$-(\gamma^{2}a_{2}+a_{1})=-\gamma(\gamma^{2}+1)+s_{2}{k}^{2}-s_{4}{k}^{4}+O({k}^{6}),$ (41)
@@
Note that the constant term $-\gamma(\gamma^{2}+1)$ in Eq. (41) is canceled by the external normal traction in Eq. (38a), and the higher-order terms of $k$ is left as the amount that the external normal traction exceeds the body force.
@@
The electrostatic contributions from the external $\rho_{m}^{2}s_{2}>0$ and the internal $\rho_{m}^{2}\frac{\gamma^{2}}{\beta}>0$ destabilize the system, serving as negative corrections to the surface tension in the linear relation with $k$. The internal contribution is known as the Lippmann tension [30]. For the cubic contribution, the external term $\rho_{m}^{2}s_{4}>0$ and the internal term $\frac{\rho_{m}^{2}\gamma^{2}}{\beta}\left(\frac{d^{2}}{12}+\frac{m_{2}}{2\gamma}\beta d\right)>0$ enhance the stabilizing effect of the bending modulus at shorter wavelength. When the conductivity is symmetric on both sides, i.e. $\gamma=1$, we can reproduce the growth rate in [20], and the electrostatic corrections to the tension and bending modulus in [18, 29]. Such linear destabilizing effect is also reported in [15] as a consequence of the surface charge discrepancy between the two lipid layers for a no-tension membrane.
The instability observed in the present system, as reported by current studies and [20], appears to be independent of ion current or transient effects. This presents a significant departure from the behavior predicted by the LDM model in [10], where, under identical conditions, the fully charged membrane remains stable. The disparity in stability predictions stemming from the same physical setup prompts the question: what underlying physical mechanisms govern this discrepancy? We will discuss the difference and elucidate the factors influencing stability in next section.
@@ IV Instantaneous charge relaxation with asymmetric membrane
In this section, we will first examine the effects of key parameters, followed by a discussion on stabilizing electrostatic body forces and destabilizing electrostatic surface traction separately. Next, we will show the spontaneous curvature induced by the asymmetric conductivity. By the end, we will compare the linear stability analysis based on LDM and EK framework.
@@
We consider the physical system with cells or giant lipid vesicles, and thus the fluids are aqueous salt solutions or water. Our analysis focuses on the case of equal-density and equal-viscosity fluids, $\rho_{1}=\rho_{2}$ and $\mu_{1}=\mu_{2}\approx 10^{-3}$ Pa s. The value of the Debye length varies from $\lambda_{D}\sim 1\;\mathrm{\mu m}$ for pure water to $\lambda_{D}\sim 10\;\mathrm{nm}$ for 1 mM NaCl. For pure water, the absolute permittivity is around $\varepsilon_{f}\approx 7\times 10^{-10}\;\mathrm{F\>m^{-1}}$. The diffusion coefficient is around $D\sim 10^{-9}\;\mathrm{m^{2}\>s^{-1}}$. The character length scale for the gap between the electrodes is $L\sim 100\;\mathrm{\mu m}$. For the lipid membrane, typical values for the thickness is around $d_{0}\sim 5\;\mathrm{nm}$, and the capacitance is around $C_{m}\sim 0.01\;\mathrm{F\;m^{-2}}$. The elastic behavior is generally characterized by the bending modulus around $\kappa\sim 10^{-20}\;\mathrm{J}$. The typical value for the stretching modulus of a lipid membrane is around $\Gamma\sim 0.2\;\mathrm{J\;m^{-2}}$ [7], and the electric field has extra effects on it, such as the compressive electric stress effectively decreasing the isotropic tension [31, 20]. In this study we focus on the case of a freely floating membrane, which is tensionless in its base state $\Gamma=0$. From these typical values we can obtain other dimensionless numbers as $Ca\sim 0.1-10$ and $\beta\sim 0.1-10$. The time scale ratio is around $\alpha\sim 0.1$, and we will consider the instantaneous charge relaxation, i.e. $\alpha=0$ in this work. The conductivity ratio $R$ can vary over a broad range [10].
IV.1 Paremeters effects
@@
Figure 3: The effect of $Ca$ on growth rate $s$ as a function of perturbation wavenumber $k$. We consider symmetric conductivity with $\gamma=1$, $V=2$, $\beta=10$ for (a) $d=0$ and (b) $d=0.1$.
@@
The growth rate in the long-wave limit Eq. (43), i.e. $k\rightarrow 0$, shows the electrostatic forces serves as the correction to the surface tension and bending modulus, which destabilize the system at the leading order, i.e. linear to $k$. The perturbation with smaller wavelength, greater $k$, will be dominated by the stabilizing effect of the elasticity. Thus, the system is unstable to a finite window of wavenumbers $0<k<k_{c}$, where $k_{c}$ is the critical wavenumber corresponding to the marginal stabile state $s=0$ as the root of Eq. (37).
@@
The effect of $Ca$ is shown in Fig. 3. The growth rate monotonically increases with $Ca$. As $Ca\rightarrow+\infty$, the elasticity effect vanishes and all wavelengths are unstable due to the electrostatic forces. In the zero-thickness limit, the growth rate at larger $k$ is dominated by the linear behavior of the internal normal traction for $Ca\rightarrow+\infty$, while for a finite thickness membrane, the the growth rate first increases then saturates to zero for $Ca\rightarrow+\infty$, which is similar to the description in LDM [10]. However, it is noteworthy that the length scale in this work is the Debye length, and thus $k\rightarrow\infty$ represents the perturbation with wavelength much smaller than Debye length, which differs from [10], where short-wave limit $k\rightarrow\infty$ is scaled with the inverse geometric length.
@@
It is interesting to notice that the effect of $\beta$ in Fig. 4(a) on the growth rate, is very similar to that of the conductivity of the membrane, $g_{m}$ in [10], Fig. 2(a), where $g_{m}=G_{m}L/\sigma$ and $G_{m}$ is the dimensional membrane conductivity. At $\beta\;(g_{m})=0$, the system is stable for all wavenumbers, while at $\beta\;(g_{m})\rightarrow+\infty$ the growth rate saturates. Mathematically, such similar behavior can be understood from the continuous current condition at the membrane for steady state in [10] (see Eq. (2.2)):
@@
which is in the same form as the Robin condition (15) at the membrane: $\bm{n}\bm{\cdot}(-\bm{E}_{i})=\beta V_{m}$, where $\beta=C_{m}\lambda_{D}/\varepsilon_{f}$. The difference on sign of the electric field, i.e. $-\bm{E}_{i}$ is due to different relative direction of surface normal $\bm{n}$ w.r.t. the potential at electrodes. Moreover, at these two limits, the internal normal stress is zero, and the system in current work has the same stress boundary at the membrane as in [10]. Thus, even if the dimensionless number $g_{m}$ and $\beta$ has very different physical meaning, and the membrane boundary condition Eq. (15) and Eq. (44) represents balance of the charge near membrane and continuity of the current separately, they do have the similar effect on the stability of the fully charged membrane mathematically.
@@
Figure 4: (a)The effect of $\beta$ on the growth rate, with black curve shows the full growth rate as in Eq. (37), while the blue curve excludes the contribution from the internal traction in Eq. (38b); (b) the effect of $\beta$ on the external and internal contribution to the growth rate in Eq. (38) at $k=1$; (c) the maximal (full) growth rate $s_{\max}$ as function of $\beta$. Note that the blue and black curve overlap at $\beta=0,\infty$ in (a). We consider the system with parameters $V=2$, $\gamma=1$, $d=0$, and $Ca=2$.

### v2

@@ III Linear stability analysis
The membrane fluctuations impose small perturbation $h(x,t)=h_{k}e^{st+ikx}$ around a flat reference state. Any variable $g$ is expanded in a series of the form $g=g^{(0)}(z,t)+g^{(1)}(z,x,t)$. The superscript $(0)$ corresponds to the base state, and $g^{(1)}=g_{k}^{(1)}e^{st+ikx}$ is the first order solution due to the effect of the small undulations. For convenience, we drop the subscript $k$ for the first order perturbation.
@@ III.1 Base state
We consider the flat membrane as the base state, $h(x)=0$, and the electric field is aligned in $z$ direction, perpendicular to the membrane. For electrostatic fields, we solve Eq. (13a) and Eq. (13b), with the Robin condition (15) and no-flux condition (7) at the membrane, and $\rho_{i}(z\rightarrow\pm\infty)=0$, $\phi_{i}(z\rightarrow\pm\infty)=\pm V$ at the far field, to obtain the charge distribution
$\rho_{1}^{(0)}=\rho_{m}e^{-\frac{z}{\gamma}},\qquad\rho_{2}^{(0)}=-\gamma^{2}\rho_{m}e^{\gamma z},$ (21)
and the base state electric potential
@@
is related to the jump in the charge density across the membrane, with $\rho^{(0)}{]\!]}_{z=0}=(1+\gamma^{2})\rho_{m}$. Note that in the base state, the current $\bm{j}_{1}^{(0)}=\bm{j}_{2}^{(0)}=\bm{0}$ for all $z$, not just at the membrane [28].
The base state electric field calculated as $\bm{E}_{i}^{(0)}=E_{z,i}^{(0)}\bm{e}_{z}=-\partial_{z}\phi_{i}^{(0)}\bm{e}_{z}$, is non-zero in the Debye layer. This represents a key difference from the base state field in LDM. In LDM, under the same physical setup, the base state field is zero unless there is ion current across the membrane [10, 7] or membrane capacitor is charging [12]. In our subsequent discussion, we will demonstrate the critical role played by the non-zero base state electric field $\bm{E}_{i}^{(0)}$ in the membrane instability. The transmembrane potential in the base state is given by
@@
where the constant internal field has been used due to the continuity of the potential at the membrane boundaries, which is consistent with $\varepsilon_{m}E_{m}^{(0)}=\varepsilon_{f}E_{z,i}^{(0)}|_{z=0}$ by substituting $\beta$ in Eq. (16). The internal field $E_{m}^{(0)}$ is obtained by keeping a finite thickness $d$, and in Sec. III.3 we will show this field introduces negative tension and results in a destabilization effect.
@@
by assuming $p_{i}^{(0)}(z\rightarrow\pm\infty)=0$. This indicates that the pressure is continuous $p^{(0)}{]\!]}_{z=0}=0$ across the flat membrane. Equation (26) also implies that in the bulk fluid there is a balance between the pressure and the Maxwell stress in the absence of viscous stress, i.e., $\nabla\cdot\bm{T}=0$.
III.2 Perturbed state equations
@@
which is solved to yield the first order potential
@@
$a_{1}=\frac{(\gamma^{2}+1)\beta+l_{2}}{l_{1}l_{2}+\beta(l_{1}+l_{2})},\quad a_{2}=\frac{(\gamma^{2}+1)\beta+\gamma^{2}l_{1}}{l_{1}l_{2}+\beta(l_{1}+l_{2})}.$ (31)
are obtained via imposing the first order of Robin condition at the membrane
$\left(h\partial_{zz}\phi_{i}^{(0)}+\partial_{z}\phi_{i}^{(1)}\right)_{z=0}=\beta{[\![}\phi^{(1)}+h\partial_{z}\phi^{(0)}{]\!]}_{z=0}.$ (32)
Note that, when subjected to perturbation, the first order current remains zero in both electrolytes $\bm{j}_{1}^{(1)}=\bm{j}_{2}^{(1)}=\bm{0}$.
@@
Rearranging Eqs. (33) gives rise to the following fourth-order differential equation for $v_{i}$:
@@
By imposing $v_{i}(z\rightarrow\pm\infty)=0$, we obtain the solution of the form $v_{i}^{(1)}=(B_{i}+C_{i}z)e^{\mp kz}$, where $B_{i}$ and $C_{i}$ are to be determined by appropriate boundary conditions. The continuity of the tangential velocity, together with the incompressibility of the interface at the first order impose $u_{1}^{(1)}|_{z=0}=u_{2}^{(1)}|_{z=0}=0$. This reduces the kinetic equation to be $v^{(1)}=\partial_{t}h=sh$, together with the continuity of the normal velocity we have $v_{1}^{(1)}|_{z=0}=v_{2}^{(1)}|_{z=0}=sh$. These boundary conditions determine the solutions as function of $s,h$. Note that, unlike the induced-charge electro-osmotic (ICEO) mechanism identified in Refs. [20, 19], the flow field in the present study is not directly driven by the electric field. Instead, it is the deformation and motion of the membrane that generates the surrounding fluid flow.
The pressure $p^{(1)}_{i}=(\partial_{z}^{3}/{k}^{2}-\partial_{z})v_{i}^{(1)}-(\rho^{(0)}\phi^{(1)})$ is obtained as
$p_{1}^{(1)}=2kshe^{-kz}-\rho_{1}^{(0)}\phi_{1}^{(1)},\qquad p_{2}^{(1)}=-2kshe^{kz}-\rho_{2}^{(0)}\phi_{2}^{(1)},$ (35)
which includes both the dynamic pressure associated with the fluid flow and electrostatic pressure due to the body force. To determine the growth rate $s$, we apply the normal traction balance at the interface
@@
The term $h\partial_{z}\left[-p^{(0)}+\frac{1}{2}(\partial_{z}\phi^{(0)})^{2}\right]$ vanishes due to the stress balance at the base state, and $\partial_{z}v^{(1)}=0$ due to the incompressibility of the membrane and the bulk fluids. From Eq. (36) we see three sources of electrostatic contribution to the membrane instability, i.e., one from the pressure difference [[p]]{[\![}-p{]\!]} due to the electrical body force, one from the external normal traction $\partial_{z}\phi^{(0)}\partial_{z}\phi^{(1)}{]\!]}$, and one from the internal normal traction $\partial_{z}\phi_{m}^{(0)}\partial_{z}\phi_{m}^{(1)}{]\!]}$ of the membrane. Note that while the base state field $\partial_{z}\phi_{m}^{(0)}$ is constant, the first order field $\partial_{z}\phi_{m}^{(1)}$ varies across the membrane (see Appendix A and Ref. [20] for detailed calculation). For a zero-thickness model, this internal traction is needed to correct the electrohydrodynamic effect as seen in Ref. [20] and Ref. [18].
@@ III.3 Linear growth rate
The growth rate is determined from Eq. (36) as
@@
where $s_{\text{ex}}$ represents the external contribution to the growth rate, including both electrical body force and normal traction from the Maxwell stress, whereas the internal contribution $s_{\text{in}}$ arises solely from normal traction:
$s_{\text{ex}}$ $=\frac{\rho_{m}^{2}}{k}[\underbrace{-(\gamma^{2}a_{2}+a_{1})}_{\text{electrical body force}}+\underbrace{\gamma(\gamma^{2}+1)}_{\text{Maxwell stress traction}}],$ (38a)
@@
The three components, along with the full growth rate in Eq. (37) are shown in Fig. 2. Note that in the zero thickness limit $d\rightarrow 0$, the internal normal traction term in Eq. (38b) reduces to
@@
contributing linearly to the growth rate. This term is in agreement with Ref. [7] in the LDM framework.
Figure 2: The components of the growth rate in Eq. (37) as a function of the perturbation wavenumber $k$, with the “total” representing the sum of the external, internal, and elastic contributions. We consider symmetric conductivity with $\gamma=1$, $V=2$, $\beta=1$, $Ca=0.2$, $\bar{\Gamma}=0.01$ and $d=0$.
Next, we examine the effect of each term in the long-wave limit $k\rightarrow 0$. With a finite thickness $d$, the internal normal traction is expanded as
@@
The electrical body force in Eq. (38a) is expanded as
$(\gamma^{2}a_{2}+a_{1})=-\gamma(\gamma^{2}+1)+s_{2}{k}^{2}-s_{4}{k}^{4}+O({k}^{6}),$ (41)
@@
Note that the constant term $\gamma(\gamma^{2}+1)$ in Eq. (41) is canceled by the external normal traction from the Maxwell stress in Eq. (38a), and the higher-order terms of $k$ represent the amount that the external normal traction from the Maxwell stress exceeds the electrical body force.
@@
The electrostatic contributions from the external term, $\rho_{m}^{2}s_{2}>0$, and the internal term, $\rho_{m}^{2}\frac{\gamma^{2}}{\beta}>0$, act to destabilize the system by effectively reducing the membrane tension. This reduction is interpreted as a negative correction to the membrane tension—an effect analogous to the classical Lippmann (electrocapillary) phenomenon [34]. For the cubic contribution, the external term $\rho_{m}^{2}s_{4}>0$ and the internal term $\frac{\rho_{m}^{2}\gamma^{2}}{\beta}\left(\frac{d^{2}}{12}+\frac{m_{2}}{2\gamma}\beta d\right)>0$ enhance the stabilizing effect of the bending modulus at shorter wavelength (greater $k$). When conductivity is symmetric on both sides, i.e. $\gamma=1$, we reproduce the growth rate in Ref. [20], and the electrostatic corrections to the tension and bending modulus in Ref. [18, 33]. Such linear destabilizing effect is also reported in Ref. [15] as a consequence of the surface charge discrepancy between the two lipid layers for a tension-free membrane.
Note that the electrostatic correction to the tension, represented by the term $\rho_{m}^{2}s_{2}$, arises from a partial cancellation between the stabilizing electrical body force and the destabilizing Maxwell stress at the membrane, as seen from Eq. (41). Specifically, the negative contribution from the body force and the positive traction due to the Maxwell stress together determine the net effect on the growth rate (38a). Since the Maxwell stress dominates in this balance, the remaining contribution $\rho_{m}^{2}s_{2}$ acts to destabilize the membrane. The physical mechanisms underlying these opposing effects are discussed in detail in Sec. IV.2.
The instability observed in the present system, as reported by this study and Ref. [20], appears to be independent of the ion current or transient effects. This presents a significant departure from the behavior predicted by the LDM model in Ref. [10], where, under identical conditions, the fully charged membrane remains stable. This difference in stability outcomes—despite originating from the same physical setting—raises an important question about the mechanisms responsible for the onset of instability. We address these differences and clarify the underlying factors governing membrane stability in the next section.
@@ IV Instantaneous charge relaxation with asymmetric membrane
In this section, we first examine the effects of key parameters, followed by a discussion of the stabilizing electrostatic body forces and destabilizing electrostatic surface traction separately. Next, we show that asymmetric electrolytes conductivities give rise to a membrane spontaneous curvature. Finally, we compare the linear stability analyses based on LDM and EK frameworks.
@@
We consider the physical parameters relevant to cells or giant lipid vesicles, and thus the fluids are aqueous salt solutions or water. Our analysis focuses on the case of equal-density and equal-viscosity fluids, $\rho_{1}=\rho_{2}$ and $\mu_{1}=\mu_{2}\approx 10^{-3}$ Pa s. The value of the Debye length varies from $\lambda_{D}\sim 1\;\mathrm{\mu m}$ for pure water to $\lambda_{D}\sim 1\;\mathrm{nm}$ for 100 mM NaCl. This range of ionic strengths reflects physiological conditions, is commonly used in biomimetic membrane experiments [22, 24, 25, 26], and enables the investigation of asymmetric ionic conductivity observed in experimental systems [23]. For pure water, the absolute permittivity is around $\varepsilon_{f}\approx 7\times 10^{-10}\;\mathrm{F\>m^{-1}}$. The diffusion coefficient is around $D\sim 10^{-9}\;\mathrm{m^{2}\>s^{-1}}$. The character length scale for the gap between the electrodes is $L\sim 100\;\mathrm{\mu m}$. For the lipid membrane, typical values for the thickness are around $d_{0}\sim 5\;\mathrm{nm}$, and the capacitance is around $C_{m}\sim 0.01\;\mathrm{F\;m^{-2}}$. The elastic behavior is generally characterized by the bending modulus around $\kappa\sim 10^{-20}\;\mathrm{J}$. The typical value for the stretching modulus of a lipid membrane is around $\Gamma\sim 10^{-5}\;\mathrm{J\;m^{-2}}$ [19], and the electric field has additional effects on it, such as the compressive electric stress effectively decreasing the isotropic tension [35, 20]. From these typical values we obtain other dimensionless numbers as $Ca\sim 0.01-10$, $\bar{\Gamma}=10^{-3}-10^{3}$ and $\beta\sim 0.01-10$. The time scale ratio is around $\alpha\sim 0.1$, and we consider the instantaneous charge relaxation, i.e. $\alpha=0$ in this work. The conductivity ratio $R$ can vary over a broad range [10].
IV.1 Effect of physical parameters
@@
Figure 3: The effect of CaCa on growth rate $s$ as a function of perturbation wavenumber $k$. We consider symmetric conductivity with $\gamma=1$, $V=2$, $\beta=1$, $\bar{\Gamma}=0.01$ for (a) $d=0$ and (b) $d=1$.
@@
The growth rate in the long-wave limit Eq. (43), i.e. $k\rightarrow 0$, shows the electrostatic forces serve as the correction to the membrane tension and bending modulus, with the tension destabilizes the system at the leading order, i.e., linear in $k$. The perturbation with smaller wavelength is dominated by the stabilizing effect of the bending elasticity. Thus, the system is unstable to a finite window of wavenumbers $0<k<k_{c}$, where $k_{c}$ is the critical wavenumber corresponding to the marginal stable state $s=0$ as the root of Eq. (37).
@@
The effect of $Ca$ is illustrated in Fig. 3. The growth rate monotonically increases with $Ca$. As $Ca\rightarrow+\infty$, the elastic effect vanishes and all wavelengths are unstable due to electrostatic forces. In the zero-thickness limit, the growth rate at larger $k$ is dominated by the linear behavior of the internal normal traction for $Ca\rightarrow+\infty$, while for a finite thickness membrane, the growth rate first increases then saturates to zero for $Ca\rightarrow+\infty$, which is similar to the description in LDM [10]. However, it is noteworthy that the length scale in this work is the Debye length, and thus $k\rightarrow\infty$ represents the perturbation with wavelength much smaller than Debye length, which differs from Ref. [10], where short-wave limit $k\rightarrow\infty$ is scaled with the inverse geometric length.
@@
It is interesting to note that the effect of $\beta$ in Fig. 4(a) on the growth rate is very similar to that of the conductivity of the membrane, $g_{m}$ in Ref. [10], Fig. 2(a), where $g_{m}=G_{m}L/\sigma$ and $G_{m}$ is the dimensional membrane conductivity. At $\beta\;(g_{m})=0$, the system is stable for all wavenumbers, while at $\beta\;(g_{m})\rightarrow+\infty$ the growth rate saturates. Mathematically, such similar behavior can be understood from the continuous current condition at the membrane for steady state in Ref. [10] (see Eq. (2.2)):
@@
which is in the same form as the Robin condition (15) at the membrane: $\bm{n}\bm{\cdot}(-\bm{E}_{i})=\beta V_{m}$, where $\beta=C_{m}\lambda_{D}/\varepsilon_{f}$. The difference in the sign of the electric field, i.e. $\bm{E}_{i}$ is due to the different relative direction of the surface normal $\bm{n}$ w.r.t. the potential at electrodes. Moreover, in these two limits, the internal normal stress is zero, and the system in the current work has the same stress boundary at the membrane as in Ref. [10]. Thus, even if the dimensionless number $g_{m}$ and $\beta$ have very different physical meanings, and the membrane boundary condition Eq. (15) and Eq. (44) represent balance of the charge near membrane and continuity of the current respectively, they do have the similar effect on the stability of the fully charged membrane mathematically.
(a)(b)(c)
@@
Figure 4: (a)The effect of $\beta$ on the growth rate, with black curve shows the full growth rate as in Eq. (37), while the blue curve excludes the contribution from the internal traction in Eq. (38b); (b) the effect of $\beta$ on the external and internal contribution to the growth rate in Eq. (38) at $k=1$; (c) the maximal (full) growth rate $s_{\max}$ as a function of $\beta$. Note that the blue and black curves overlap at $\beta=0,\infty$ in (a). We consider the system with the parameters $V=3$, $\gamma=1$, $d=0$, $Ca=0.5$, and $\bar{\Gamma}=0.01$.